### v1

# Gamma: Toward Generic Image Assessment with Mixture of Assessment Experts

Abstract
Image assessment aims to evaluate the quality and aesthetics of images and has been applied across various scenarios, such as natural and AIGC scenes. Existing methods mostly address these sub-tasks or scenes individually. While some works attempt to develop unified image assessment models, they have struggled to achieve satisfactory performance or cover a broad spectrum of assessment scenarios. In this paper, we present Gamma, a Generic imAge assessMent model using Mixture of Assessment Experts, which can effectively assess images from diverse scenes through mixed-dataset training. Achieving unified training in image assessment presents significant challenges due to annotation biases across different datasets. To address this issue, we first propose a Mixture of Assessment Experts (MoAE) module, which employs shared and adaptive experts to dynamically learn common and specific knowledge for different datasets, respectively. In addition, we introduce a Scene-based Differential Prompt (SDP) strategy, which uses scene-specific prompts to provide prior knowledge and guidance during the learning process, further boosting adaptation for various scenes. Our Gamma model is trained and evaluated on 12 datasets spanning 6 image assessment scenarios. Extensive experiments show that our unified Gamma outperforms other state-of-the-art mixed-training methods by significant margins while covering more scenes.
Machine Learning, ICML
1 Introduction
Image assessment is a long-standing research topic in the field of image processing, primarily comprising two tasks: Image Quality Assessment (IQA) and Image Aesthetic Assessment (IAA). These tasks require models to automatically evaluate the visual quality and aesthetic appeal of images, respectively. They have broad applications in various real-world scenarios, such as guiding image dehazing (Zhao et al., 2021), selecting high-quality images in a data engine (Rombach et al., 2022), serving as tools in an agentic system (Yang et al., 2024), or acting as reward models when aligning image generative models with human feedback (Liang et al., 2024).
Due to differences in image content and application scenarios, image assessment has spawned many sub-tasks, such as Natural-IQA for natural images, Face-IQA for facial images, AIGC-IQA for generated images, and IAA. Accordingly, numerous methods (Ke et al., 2021; He et al., 2022; Su et al., 2023b) have been proposed to address these specific tasks. However, these models often struggle to apply directly to other scenes or typically require task-specific fine-tuning on a given dataset. As illustrated in Figure 1, it is challenging for DEIQT (Qin et al., 2023) to transfer to other datasets without fine-tuning. This limitation prevents image assessment models from being widely applicable, e.g., IQA models are needed to assess facial, artistic, and natural images in the AIGC scene. Hence, there is an urgent need for a model that can effectively handle a variety of scenarios.
Figure 1: “DEIQT Finetune” is trained and tested on the same dataset. “DEIQT ZS” directly assesses images using the model trained on other datasets, which perform poorly. PromptIQA and our Gamma are trained on mixed datasets. Our Gamma performs well on multiple datasets simultaneously and even surpass the task-specific method DETQT.
Figure 2: Images with similar MOS labels from different datasets exhibit drastically different perceptual quality. It is not hard to observe that image (a) has clearly superior quality than the other three. Zoom in for a better view.
To mitigate this issue, some approaches attempt to combine many datasets from various assessment tasks to train a general image assessment model. For instance, UNIQUE (Zhang et al., 2021) and LIQE (Zhang et al., 2023) utilize multiple authentic and synthetic natural IQA datasets for mixed training, but they focus only on Natural-IQA. Q-Align (Wu et al., 2023) uses a large language model with billions of parameters to unify IQA and IAA tasks, but it has a low inference speed and focuses solely on natural images. Additionally, PromptIQA (Chen et al., 2024b) employs image-score pairs as prompts for quality predictions, but it is inflexible as it requires multiple images as references during inference. These methods, however, fail to achieve competitive performance compared to models trained on specific datasets and cover a broad range of scenes. Thus, it is imperative to develop a foundational image assessment model capable of evaluating images from various scenes.
To this end, we present Gamma, a Generic imAge assessMent model using Mixture of Assessment experts, to achieve unified image assessment across multiple datasets effectively. We found that the primary challenge in mixed-dataset training is the mean opinion score (MOS) bias between different datasets, e.g., images with similar MOS may exhibit different perceptual qualities across various datasets. As shown in Figure 2, samples from KonIQ (Hosu et al., 2020), UWIQA (Yang et al., 2021a), and SPAQ (Fang et al., 2020) show obvious differences in perceptual quality despite being labeled with similar MOS. To address this challenge, we introduce a novel Mixture of Assessment Experts (MoAE) module in our Gamma model. MoAE consists of two types of experts: shared experts and adaptive experts. The shared experts are employed throughout to learn dataset-shared knowledge, while the adaptive experts are dynamically activated to varying degrees to learn dataset-specific knowledge. Additionally, an image-based router modulates the contributions of each adaptive expert. This strategy allows the model to capture common features while flexibly adjusting representative features for different datasets. Compared with general Mixture of Experts (MoE) (Shazeer et al., 2017) and Lora-based MoE (Liu et al., 2024), we equip the MoAE module only in the rear blocks instead of all blocks, making it more efficient. Secondly, we propose a Scene-based Differential Prompt (SDP), which uses different prompts for different datasets according to their scenes. This strategy provides scene-specific knowledge for representation learning of different datasets, guiding the mixed-dataset training process.
Our Gamma model is uniformly trained on a mixture of 12 datasets from 6 image assessment scenarios spanning IQA and IAA tasks. We then evaluate it on 12 datasets, demonstrating that it not only outperforms state-of-the-art (SOTA) mixed-training methods by notable margins, but also covers more scenarios. In some benchmarks, Gamma even surpasses some SOTA task-specific methods. Additionally, if we fine-tune our MoAE-equipped Gamma on specific datasets, it can achieve SOTA performance on 12 datasets. Moreover, the unified pre-trained Gamma can be utilized as a foundational model to significantly enhance other image assessment tasks, e.g., medical image quality assessment, and can achieve SOTA performance after task-specific training. In summary, our contributions are:
  • We present Gamma, a powerful and generic image assessment model, capable of accurately assessing images from various scenarios through mixed training.
  • We propose a novel Mixture of Assessment Experts (MoAE) module to extract representative features adaptively and a Scene-based Differential Prompt (SDP) strategy to provide guidance for representation learning, thereby achieving effective mixed-dataset training.
  • Extensive experiments show that Gamma achieves SOTA performance on 12 datasets across 6 image assessment scenes in both mixed training and task-specific training settings.
2 Related Work
2.1 Image Assessment
Image Assessment mainly includes two subtasks: Image Quality Assessment (IQA) and Image Aesthetic Assessment (IAA). In the deep learning era, these two tasks have achieved significant breakthroughs. For the IQA task, researchers develop various advanced techniques to improve performance, including multi-level feature aggregation (Li et al., 2018; Chen et al., 2024a; Xu et al., 2024; Zhang et al., 2018; Mittal et al., 2012b; Ying et al., 2020), adaptive convolution (Su et al., 2020), transformer methods (Ke et al., 2021; Qin et al., 2023), vision-language models (VLMs) (Wang et al., 2023; Zhang et al., 2023) and large language models (LLM) (You et al., 2023). Moreover, besides the natural image assessment, these are various IQA methods for other scenes, such as face IQA (Ou et al., 2021; Su et al., 2023b; Jo et al., 2023), AIGC IQA (Yuan et al., 2023), underwater IQA (Yang et al., 2021b; Guo et al., 2023; Yang & Sowmya, 2015; Liu et al., 2023). For the IAA task, numerous methods have also been proposed to improve performance, including loss function (Talebi & Milanfar, 2018), novel transformer architecture (Tu et al., 2022), multi-level features (Hosu et al., 2019), theme information (He et al., 2022; Li et al., 2023b) and multimodal pre-training (Ke et al., 2023).
2.2 Mixed Training for Image Assessment
As a fundamental image processing task, image assessment has been effectively applied to various scenarios. Recently, some works have attempted to develop unified methods that can be used in multiple IQA settings through mixed-dataset training. UNIQUE (Zhang et al., 2021) sampled pairs of images from IQA datasets and computes the probability that the first image of each pair is of higher quality. StairIQA (Sun et al., 2023) designed separate IQA regression heads for each dataset. PromptIQA (Chen et al., 2024b) utilized a short sequence of Image-Score Pairs as prompts for quality predictions. Q-Align (Wu et al., 2023) used large language model (LLM) to unify IQA and IAA tasks. However, most existing works fail to achieve competitive performance with task-specific methods and do not cover various scenes. Our Gamma outperforms other SOTA mixed-training methods by significant margins while covering more scenes.
2.3 Mixture of Experts
Mixture-of-Experts (MoE) divides specific parts of the parameters into several subsets, each of which is called an expert. It sets up a router that assigns experts to different inputs. Recently, the MoE structure has achieved remarkable success in large language models (LLM). For instance, DeepSeekMoE (Dai et al., 2024) proposes a novel MoE architecture that uses shared and routed experts to extract common and dynamic knowledge simultaneously. Beyond the natural language processing tasks, the idea of MoE has also been applied to vision models (Dai et al., 2021; Riquelme et al., 2021; Chen et al., 2023) and multimodal transformers (Wang et al., 2022; Feng et al., 2023). In Gamma, we utilize novel MoE module to effectively learn specific and general features of multi-dataset.
Figure 3: (a) The architecture of Gamma: It consists of a visual encoder $\mathcal{V}$ and text encoder $\mathcal{T}$; We add the Mixture of Assessment Experts (MoAE) to the last few layers of both encoders. We introduce a Scene-based Differential Prompt (SDP) to prompt images from different scenes (See Section 3.4 for details). (b) The MoAE module: It involves one shared expert ($E^{\mathrm{shared}}$) and several adaptive experts (from $E_{1}^{\mathrm{adaptive}}$ to $E_{n}^{\mathrm{adaptive}}$). We employ a router to adaptively activate the adaptive experts. We then use a learnable factor $\sigma$ to merge the features of two type of experts.
3 Method
3.1 Preliminary
As a foundational vision-language model (VLM), CLIP (Radford et al., 2021) has shown significant promise in supporting a wide array of vision tasks. Specifically, CLIP is composed of a transformer-based visual encoder $\mathcal{V}$ and a text encoder $\mathcal{T}$, which generate aligned visual representations $I$ and text representations $T$ for each image-text pair. Utilizing these features, we can compute cosine similarity scores between image and text pairs across different domains or tasks to perform task-specific predictions, including image assessment tasks. Recently, to enhance CLIP’s capabilities in the field of image assessment, UniQA (Zhou et al., 2024b) fine-tuned CLIP on large-scale synthetic and authentic image-text datasets focused on image quality and aesthetics. This approach demonstrates excellent performance on both IQA and IAA tasks after task-specific fine-tuning. However, the model lacks foundational applicability across various image assessment scenarios without fine-tuning. Building on the unified training pipeline proposed in UniQA, we propose two approaches to confront MOS bias present in different datasets and develop a foundational image assessment model. In the following, we will provide a detailed exposition of its components.
3.2 Overview of Gamma
As illustrated in Figure 3, our Gamma employs a visual encoder $\mathcal{V}$ to extract visual features $I\in\mathbb{R}^{d}$, and a text encoder $\mathcal{T}$ to extract text features $T\in\mathbb{R}^{d}$. After these encoders, a tunable adapter is used to obtain a score $\mathrm{q}$ representing image quality or aesthetics, following the methods in (Zhang et al., 2023) and (Zhou et al., 2024b). This process can be described as:
$\mathrm{q}=\sum_{k=1}^{5}\mathrm{C}_{k}\mathrm{Softmax}({I^{\prime}}^{\top}T_{k}/\tau),I^{\prime}=\mathrm{Adapter}(I),$ (1)
[…]
3.3 Mixture of Assessment Experts
To develop a unified and generic image assessment model, we aim to combine multiple image assessment datasets for joint training. Unfortunately, the mean opinion score (MOS) introduces significant biases across different datasets, which hinders joint training. To address this challenge, we propose the MoAE module, where several experts are employed to learn the diverse biases of different datasets. As shown in Figure 3 (b), the proposed MoAE module includes a shared assessment expert ($E^{\mathrm{shared}}$) to learn common knowledge of image assessment and several adaptive assessment experts (from $E_{1}^{\mathrm{adaptive}}$ to $E_{n}^{\mathrm{adaptive}}$) to dynamically learn dataset-specific knowledge.
The Shared Assessment Expert. The shared assessment expert $E^{\mathrm{shared}}$ inherits the image assessment capabilities of the original CLIP model by reusing its weights. This expert remains frozen during training to ensure that the learned world knowledge is retained. Thus, the model can capture common representations across various contexts and maintain its original multi-modal capabilities. Given an input hidden state $x\in\mathbb{R}^{d}$, the output of the shared assessment expert is:
$y^{\mathrm{shared}}=E^{\mathrm{shared}}(x),$ (2)
where $E^{\mathrm{shared}}(\cdot)$ is implemented as the original feed-forward network (FFN) of the CLIP model.
Figure 4: Visual examples from different datasets, which include natural images, underwater images, face images, AI-generated images, and etc. The first value is the prediction score and the second value is the ground-truth MOS. Our Gamma can accurately evaluate images from different scenes, demonstrating the generalization and effectiveness. All images are resized for better visibility.
The Adaptive Assessment Expert. The adaptive assessment expert module contains two components: (1) $n$ experts $\{E_{i}^{\mathrm{adaptive}}\}_{i=1}^{n}$ to capture diverse facets of multi-dataset information; and (2) a router $G(\cdot)$ to tailor the contribution of different experts based on the input feature. Given an input feature $x\in\mathbb{R}^{d}$, the output $y^{\mathrm{adaptive}}$ can be computed as:
$y^{\mathrm{adaptive}}$ $=\sum_{i=1}^{n}G(x)_{i}E_{i}^{\mathrm{adaptive}}(x),$ (3)
$G(x)$ $=\mathrm{Softmax}(Wx).$
Here, the router $G(\cdot)$ is a linear transformation for the input feature $x$; $W\in\mathbb{R}^{n\times d}$ is the transformation matrix. To avoid unreasonable weights, we utilize a $\mathrm{Softmax}$ operator to normalize the contribution weights. This ensures that the model can learn dataset-specific knowledge efficiently. Notably, we initialize these experts with the weights of pre-trained CLIP to equip them with image assessment capabilities.
MoAE Module. Based on the above experts, the MoAE module merges the features of the two types of experts with a learnable factor $\sigma$, as shown on the right side of Figure 3. Thus, the output of the MoAE module can be expressed as:
$y^{\mathrm{MoAE}}=y^{\mathrm{shared}}+\sigma\cdot y^{\mathrm{adaptive}}.\$ (4)
The $\sigma$ factor is zero-initialized so that the visual and text encoders can generate aligned features at the beginning. In practice, we freeze the shared experts and set the adaptive experts to be tunable only. This approach maintains parameter efficiency during mixed training and preserves the multi-modal capabilities of the original model.
We incorporate the MoAE module into the last $K$ layers of both visual and text encoders, as shown in Figure 3. This strategy makes our method both effective and efficient. The visual feature extraction process can be formulated as:
$I_{i}$ $=\bm{\mathcal{V}}_{i}(I_{i-1}),\quad i=1,2,\ldots,L-K$ (5)
$I_{j}$ $=\bm{\mathcal{V}}_{j}^{\mathrm{MoAE}}(I_{j-1}),\quad j=L-K+1,\ldots,L$
where $L$ denotes the number of layers of the visual encoder; $I_{i}$ represents the visual features of the $i$-th encoder layer; and $\bm{\mathcal{V}}^{\mathrm{MoAE}}$ represents the MoAE-equipped visual encoder layer. The text branch operates similarly to the visual branch.
3.4 Scene-based differential prompt
To facilitate scene-guided learning, we introduce a Scene-based Differential Prompt (SDP) to help the model acquire scene-specific knowledge from different datasets. We utilize 12 datasets spanning 6 image assessment scenarios, including synthetic distortion nature IQA, authentic distortion nature IQA, face IQA, AIGC IQA, underwater IQA, and IAA, for mixed training. We categorize these datasets into five groups based on their scenes: natural quality, AI-generated quality, underwater quality, face quality, and natural aesthetics. Specifically, for the face quality assessment dataset, we use prompts such as face bad-quality, face poor-quality, face fair-quality, face good-quality, face perfect-quality appended to the word image. For more details on these prompts, please refer to Appendix A.4.
This strategy effectively differentiates the feature space of images from various scenes and enhances scene-specific knowledge, thereby mitigating the MOS bias across different datasets. Experimental results show that this method significantly improves the model’s performance (Table 1).
4 Experiments
4.1 Datasets
We utilize 12 datasets for unified training and testing, encompassing both image quality and aesthetic assessment tasks. For the IQA task, five different assessment scenarios are included: synthetic distortion nature IQA (SDN-IQA), authentic distortion nature IQA (ADN-IQA), face IQA (F-IQA), AIGC IQA (AG-IQA), and underwater IQA (U-IQA). For the IAA task, we use two classical benchmarks, AVA and AADB. In addition, we use two rare datasets to verify the generalization ability of the model, i.e., exBeDDE and ECIQAD. The exBeDDE is a dehazed image quality assessment (D-IQA) dataset, and the ECIQAD is an enhanced colonoscopy image quality assessment (EC-IQA) dataset. Details about these datasets is provided in Table 9.
Table 1: Comparison with SOTA task-specific and mixed-training models on 12 datasets for 6 image assessment tasks. “Gamma” and “Gamma-T” denote the mixed-training and task-specific models, respectively. Gamma uses the Scene-based Differential Prompt (SDP) for training and testing. indicates that we retrain the model with the same data split as ours.
Task Synthetic Distortion Nature IQA (SDN-IQA) Authentic Distortion Nature IQA (ADN-IQA)
Dataset LIVE CSIQ TID2013 KADID LIVEC KonIQ SPAQ
Training Type Method SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
HyperIQA 0.962 0.966 0.923 0.942 0.840 0.858 0.852 0.845 0.859 0.882 0.906 0.917 0.911 0.915
MUSIQ 0.940 0.911 0.871 0.893 0.773 0.815 0.875 0.872 0.702 0.746 0.916 0.928 0.918 0.921
TOPIQ 0.943 0.942 0.908 0.925 0.813 0.845 0.877 0.875 0.833 0.868 0.915 0.925 0.914 0.917
Task Specific DEIQT 0.980 0.982 0.946 0.963 0.892 0.908 0.889 0.887 0.875 0.894 0.921 0.934 0.919 0.923
LoDA 0.975 0.979 - - 0.869 0.901 0.931 0.936 0.876 0.899 0.932 0.944 0.925 0.928
UniQA 0.981 0.983 0.963 0.973 0.916 0.931 0.940 0.943 0.890 0.905 0.933 0.941 0.924 0.928
Gamma-T 0.982 0.971 0.973 0.978 0.944 0.950 0.960 0.961 0.899 0.921 0.945 0.952 0.928 0.931
UNIQUE 0.969 0.968 0.902 0.927 - - 0.878 0.876 0.854 0.890 0.896 0.901 - -
LIQE 0.972 0.953 0.946 0.943 - - 0.932 0.933 0.902 0.908 0.920 0.905 - -
StairIQA 0.937 0.934 0.768 0.843 0.675 0.773 0.785 0.805 0.780 0.855 0.865 0.896 0.903 0.907
Mixed Training PromptIQA 0.936 0.934 0.926 0.939 0.903 0.922 0.928 0.931 0.913 0.928 0.929 0.943 0.923 0.926
Gamma 0.957 0.952 0.949 0.966 0.926 0.934 0.960 0.962 0.851 0.871 0.940 0.949 0.923 0.928
Gamma 0.953 0.953 0.960 0.968 0.935 0.941 0.962 0.964 0.891 0.914 0.939 0.950 0.929 0.932
Task Face IQA (F-IQA) AIGC IQA (AG-IQA) Underwater IQA (U-IQA) Image Aesthetic Assessment (IAA)
Dataset GFIQA20k Dataset AGIQA3k Dataset UWIQA Dataset AVA Dataset AADB
Training Type Method SRCC PLCC Method SRCC PLCC Method SRCC PLCC Method SRCC PLCC Method SRCC PLCC
SDD-FIQA 0.602 0.649 DBCNN 0.821 0.876 FDUM 0.694 0.689 MaxViT 0.708 0.745 MUSIQ 0.706 0.712
Task Specific IFQA 0.697 0.722 HyperNet 0.836 0.890 UCIQE 0.627 0.626 TANet 0.758 0.765 TANet 0.738 0.737
TOPIQ 0.966 0.967 CLIPIQA 0.843 0.805 URanker 0.674 0.663 VILA 0.774 0.774 TAVAR 0.761 0.763
GPFIQA 0.964 0.965 PSCR 0.850 0.906 UIQI 0.742 0.741 UniQA 0.776 0.776 UniQA 0.786 0.787
Gamma-T 0.968 0.968 Ours-T 0.894 0.921 Ours-T 0.870 0.880 Ours-T 0.785 0.784 Ours-T 0.793 0.798
UNIQUE - - UNIQUE - - UNIQUE - - UNIQUE - - UNIQUE - -
LIQE - - LIQE - - LIQE - - LIQE - - LIQE - -
StairIQA 0.937 0.935 StairIQA 0.755 0.833 StairIQA 0.722 0.727 StairIQA - - StairIQA - -
Mixed Training PromptIQA 0.970 0.971 PromptIQA 0.851 0.901 PromptIQA 0.877 0.884 PromptIQA - - PromptIQA - -
Gamma 0.970 0.970 Gamma 0.870 0.910 Gamma 0.863 0.878 Gamma 0.740 0.737 Gamma 0.742 0.743
Gamma 0.970 0.970 Gamma 0.887 0.923 Gamma 0.873 0.884 Gamma 0.750 0.749 Gamma 0.756 0.755
Table 2: The effect of Mixture of Assessment Experts (MoAE) and Scene-based Differential Prompt (SDP).
Dataset LIVEC KonIQ LIVE CSIQ AGIQA3k UWIQA AVA
MoAE SDP SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
$\times$ $\times$ 0.765 0.792 0.858 0.885 0.927 0.918 0.852 0.898 0.800 0.866 0.750 0.768 0.681 0.672
$\times$ $\checkmark$ 0.843 0.856 0.874 0.896 0.929 0.917 0.866 0.901 0.841 0.887 0.770 0.780 0.721 0.715
$\checkmark$ $\times$ 0.851 0.871 0.940 0.949 0.957 0.952 0.949 0.966 0.870 0.910 0.863 0.878 0.740 0.737
$\checkmark$ $\checkmark$ 0.891 0.914 0.939 0.950 0.960 0.968 0.953 0.953 0.887 0.923 0.873 0.884 0.750 0.749
Table 3: The impact of different number of experts in adaptive experts. We use naive prompt strategy for ablation.
Dataset LIVEC CSIQ TID2013 AGIQA3k UWIQA
Experts Number SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
Zero Expert 0.765 0.792 0.852 0.898 0.792 0.826 0.800 0.866 0.750 0.768
One Expert 0.842 0.866 0.945 0.963 0.918 0.931 0.866 0.908 0.859 0.873
Three Experts 0.851 0.871 0.949 0.966 0.926 0.934 0.870 0.910 0.863 0.878
Five Experts 0.854 0.889 0.951 0.965 0.926 0.934 0.872 0.911 0.860 0.876
4.2 Implementation Details
Following the settings in (Su et al., 2020; Ke et al., 2021), we randomly divide each dataset into 80% for training and 20% for testing. The training dataset is a mixture of the training sets of each dataset and we test Gamma on each test data separately. This process is repeated 10 times, and the median of the 10 scores is reported as the final score. We use the pre-trained weight of UniQA (Zhou et al., 2024b), which uses CLIP-B/16 as multimodal encoder. We freeze the CLIP visual and text encoders, training only the MoAE module and adapter. For the unified training, we train the model for 10 epochs with a batch size of 8. The initial learning rate is set to 2e-5. We normalize the MOS/DMOS scale to [0, 1] and utilize Adam optimizer (Kingma, 2014) and MSE loss to optimize the model. For the task-specific training, we use different training settings according to the task and size of datasets. More training details are provided in the appendix.
Table 4: The impact of different model configuration in the proposed MoAE module.
Dataset LIVEC CSIQ AGIQA3k UWIQA AVA
Model Configuration SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
Unfreeze shared expert 0.849 0.869 0.946 0.957 0.864 0.904 0.858 0.871 0.720 0.719
w/o Merging factor $\sigma$ 0.847 0.866 0.932 0.947 0.851 0.903 0.845 0.869 0.698 0.697
Our MoAE module 0.851 0.871 0.949 0.966 0.870 0.910 0.863 0.878 0.740 0.737
Table 5: The impact of adding MoAE to different numbers of layers for training. Time means training time.
Dataset Parms Time LIVEC KonIQ LIVE CSIQ AGIQA3k UWIQA AVA
MoE Layer Million Hours SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
w/o MoAE 149.9 3.5 0.765 0.792 0.858 0.885 0.927 0.918 0.852 0.898 0.800 0.866 0.750 0.768 0.681 0.672
Last 4 layers 231.8 7.5 0.830 0.859 0.933 0.944 0.954 0.952 0.937 0.960 0.866 0.909 0.853 0.867 0.735 0.732
Last 6 layers 272.7 10.2 0.851 0.871 0.940 0.949 0.957 0.952 0.949 0.966 0.870 0.910 0.863 0.878 0.740 0.737
Last 8 layers 313.6 13.4 0.852 0.883 0.941 0.947 0.956 0.951 0.953 0.967 0.872 0.913 0.866 0.875 0.746 0.743
All 12 layers 395.5 17.2 0.860 0.883 0.939 0.950 0.954 0.950 0.954 0.968 0.881 0.908 0.863 0.868 0.728 0.725
4.3 Main Results
Our MoAE-equipped model can be used for task-specific training and mixed training, both of which can achieve state-of-the-art (SOTA) performance, as shown in Table 1.
Task-specific Training. We apply our method to 12 image assessment datasets. We use the fixed naive prompt (described in Section 3.2) for training and testing. We observe that our method outperforms all others methods by a significant margin. On some benchmark, our method achieve dramatic improvements, such as SRCC of 0.944 (v.s. 0.916) on TID2013 and 0.945 (v.s. 0.933) on KonIQ. Since images in these 12 datasets encompass a wide variety of contents and distortion types, it is particularly challenging to consistently achieve the leading performance on all of them.
Mixed Training. We conduct mixed training on 12 image assessment datasets. The trained model can be used to assess the images from these datasets. The experimental results are reported in Table 1. When compared with other mixed training models, such as StairIQA and PromptIQA, our method exhibits powerful and superior performance on each dataset. More importantly, our method can also be used to IAA tasks and demonstrates excellent performance. It is worth noting that our mixed training model even achieves results comparable to task-specific models on datasets such as KADID, KonIQA, SPAQ, and GFIQA. These results demonstrate that our approach can be effectively applied to different image assessment scenarios.
Comparison with other mixed training methods. We conduct a detailed comparison with other mixed training methods, including LIQE, UNIQUE and Q-Align, with the same training datasets. We also perform the cross-dataset validation with LIQE and UNIQUE. Please refer to Appendix B for datails.
Qualitative Results. We visualize the image assessment results from different datasets, covering various scenarios, as shown in Figure 4. We can notice that our Gamma can accurately assess images from various tasks. These results shows the high generalization capability of our Gamma.
4.4 Ablation Studies
We conduct detailed ablation studies to validate the effectiveness of our proposed modules. Note that we use naive prompt strategy (described in Section 3.2) to perform all ablations unless otherwise specified. We uniformly use 12 datasets for ablation experiments. Considering the page limit, we only show the datasets with relatively large differences in results. More ablations about prompt sensitivity and model efficiency please refer to Appendix C and D.
Effectiveness of the prompt strategy. We propose the Scene-based Differential Prompt (SDP) to prompt different datasets. We evaluate the effectiveness of this strategy in Table 1. We can notice that the SDP strategy can improve the model performance on multiple datasets, especially on CSIQ (+1.1% SRCC), LIVEC (+4% SRCC) and AGIQA-3k (+1.7 % SRCC). These results demonstrate that the SDP can effectively guide model learn differential features for different datasets, thus enhancing model performance. Furthermore, we ablate the SDP strategy and MOAE module respectively to explore their relationship and impact on model performance. As shown in Table 2, both methods can improve the performance of the model, such as +7.8% SRCC of SDP and +8.6% SRCC of MoAE on LIVEC. This shows the effectiveness of this adaptive expert feature learning and text guidance for multi-dataset learning. When the two methods are used simultaneously, the model can achieve the best results. Therefore, the two methods are mutually beneficial.
Figure 5: (a) Average performance when using different numbers of experts. Adding adaptive experts significantly improves model performance. (b) Average activations of three experts in the last layer of the visual encoder with naive prompts. Image evaluations of different scenes have different activation patterns.
The number of experts. We explore the impact of different numbers of experts in the adaptive assessment experts. As shown in Table 3 and Figure LABEL:fig:ab-expert, the model achieves higher performance with more experts. This suggests that adding experts can better cope with the dataset bias problem when using a mixed training strategy. Using five experts cannot significantly improve the performance as the performance approaches saturation. We use three experts to constitute the adaptive assessment experts in MoAE to achieve the optimal trade-off between accuracy and efficiency.
Table 6: Results when only one adaptive expert is activated. The weights factors of other experts are set to 0. It can be observed that different experts focus on different datasets.
Dataset LIVEC KonIQ LIVE CSIQ AGIQA3k UWIQA GFIQA AVA
Expert index SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
$1$-th expert 0.847 0.860 0.927 0.938 0.933 0.933 0.894 0.906 0.815 0.870 0.770 0.779 0.959 0.957 0.666 0.673
$2$-th expert 0.715 0.672 0.681 0.717 0.900 0.861 0.815 0.846 0.832 0.885 0.755 0.756 0.826 0.797 0.663 0.652
$3$-th expert 0.768 0.741 0.794 0.818 0.918 0.917 0.833 0.877 0.808 0.910 0.691 0.709 0.903 0.897 0.715 0.716
Gamma 0.851 0.871 0.940 0.949 0.957 0.952 0.949 0.966 0.870 0.910 0.863 0.878 0.970 0.970 0.740 0.737
Table 7: Generalization capability validation on the exBeDDE and ECIQAD datasets. The “Pretrained weight” denotes the model weight of mixed training. We can notice that loading pretrained weight for initialization can improve model performance.
(a) Results on the exBeDDE datasets.
Method SRCC PLCC
BRISQUE (Mittal et al., 2012a) 0.890 0.906
PSQA-I (Liu et al., 2019) 0.907 0.924
HyperIQA (Su et al., 2020) 0.917 0.926
FADE (Choi et al., 2015) 0.714 0.729
DHQI (Min et al., 2018) 0.919 0.939
VDA-DQA (Guan et al., 2022) 0.923 0.942
Ours 0.916 0.938
Ours + Pretrained weight 0.937 0.951
(b) Results on the ECIQAD datasets.
Method SRCC PLCC
BRISQUE (Mittal et al., 2012a) 0.436 0.459
BIQME (Gu et al., 2017) 0.770 0.768
BPRI (Min et al., 2017) 0.152 0.181
FRIQUEE (Ghadiyaram & Bovik, 2017) 0.663 0.656
CIQA (Chen et al., 2021) 0.738 0.735
ECIQ (Ke et al., 2021) 0.839 0.842
Ours 0.912 0.922
Ours + Pretrained weight 0.917 0.927
Effect of freezing the shared expert. We freeze the shared expert in the MoAE to retrain the multimodal capability of original model. This strategy also helps model capture the generalizable and common representation across varying contexts. Table 4 validates this method and shows that it is effective across various datasets.
Merging features with factor $\sigma$. Table 4 also demonstrates the effect of merging features of shared and adaptive experts with factor $\sigma$. We notice that this strategy improves the model performance on different datasets, especially on AVA and AGIQA3k. These results show that it is beneficial to utilize aligned features at the beginning of training and partially using features from adaptive experts.
Adding adapter to last few layers. We add the proposed MoAE module into the last few layers of the visual and text encoders. We compare the performance of adding MoAE to different numbers of layers in Table 5. After using MoAE, the performance of the model is significantly improved. We also observe that adding more than six layers of adapters does not improve model performance significantly, but further increases the model parameter and training overhead. Therefore, we integrate MoAE module in the last six layers of both visual and text encoder.
Activation patterns of different datasets. We visualize the average activation degree of three experts in the last layer of Gamma’s visual encoder for different datasets, as shown in Figure LABEL:fig:ab-activation. We can observe that the activation patterns are different for different scenarios. Specifically, the natural image assessment datasets, e.g., LIVE, CSIQ, KADID, show different activation patterns from the face IQA dataset GFIQA and the underwater IQA dataset UWIQA. The synthetic distortion and authentic distortion dataset in nature IQA also have different activation patterns. These indicate that our MoAE module can assign experts with different activation levels to images of different scenarios, thereby capturing the discriminative features effectively.
Analysis of the adaptive experts. We add an experiment in which we only use one adaptive expert and set the router weights of the other experts to 0, to explore the preferences of different experts for different datasets. As shown in Table 6, the first expert performs well on most datasets, indicating it learns a general image assessment ability. The second and third experts focus on AIGC IQA and IAA tasks, respectively, and the third expert also shows excellent evaluation capabilities for natural images. These results indicate that different experts have learned domain-specific features of different datasets. They collaborate to achieve the powerful image assessment model Gamma.
4.5 Generalization Capability Validation
We further validate the generalization capability of our method on two datasets, exBeDDE and ECIQAD. The exBeDDE is a dehazed IQA dataset and the ECIQAD is an enhanced colonoscopy IQA dataset, which belong to completely different evaluation domains compared to the used datasets in mixed training. We use naive prompt strategy for training and testing. The experimental results are reported in Table 7. We notice that our method can achieve competitive performance on these two datasets, showing the effectiveness and generalization capability of our method. More importantly, when we load the pretrained weight of Gamma for initialization, the performance of both datasets is improved and our method achieves the SOTA results. This indicates that our pretrained Gamma can be an effective foundation model to aid other assessment fields.
5 Conclusion
This paper introduces Gamma, a generic image assessment model that can be applied to various image scenarios. To achieve this, we utilize the mixed training of different datasets to obtain the assessment abilities of different scenarios. We propose a Mixture of Assessment Expert (MoAE) module and a Scene-based Differential Prompt (SDP) strategy to effectively cope with the MOS bias in different datasets. MoAE utilizes shared experts and adaptive experts to extract common and representative features adaptively. SDP strategy employs different prompts for different datasets to provide guidance for feature learning. Extensive experiments demonstrate that our method can achieve SOTA performance on various datasets simultaneously, showing the strong generalization and general image assessment capabilities.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A More Implementation Details
A.1 Evaluation Criteria
We use Spearman’s Rank-Order Correlation Coefficient (SRCC) and Pearson’s Linear Correlation Coefficient (PLCC) as criteria to measure the performance of IQA models. Both coefficients range from 0 to 1, with higher values indicating better performance.
A.2 Training Details
We follow the typical training strategy to fine-tune each dataset, including random cropping and random horizontal flipping. We conduct all experiments on 3090 GPU. Mixed training of the 12 datasets takes 10 hours on a 3090 GPU. For the task-specific training, Table 8 shows the detailed training setting for the different datasets. We use the learning rate of 2e-5 for all datasets.
Table 8: Training settings for different datasets.
Dataset Task Epoch Batch size
LIVE (Sheikh et al., 2006) SDN-IQA 50 8
CSIQ (Larson & Chandler, 2010) SDN-IQA 50 8
TID2013 (Ponomarenko et al., 2013) SDN-IQA 20 8
KADID (Lin et al., 2019) SDN-IQA 20 8
CLIVE (Ghadiyaram & Bovik, 2015) ADN-IQA 50 8
KonIQ (Hosu et al., 2020) ADN-IQA 20 8
SPAQ (Fang et al., 2020) ADN-IQA 20 8
GFIQA20k (Su et al., 2023b) F-IQA 10 8
AGIQA3k (Li et al., 2023a) AG-IQA 20 8
UWIQA (Yang et al., 2021a) U-IQA 50 8
AVA (Murray et al., 2012) IAA 20 128
AADB (Kong et al., 2016) IAA 20 8
exBeDDE (Zhao et al., 2020) D-IQA 20 8
ECIQAD (Yue et al., 2023) EC-IQA 20 8
A.3 Datasets
In this paper, we use a total of 14 datasets, 12 of which are used for unified training and 2 are used to evaluate the generalization ability of our model. We present the details of the used datasets in Table 9.
Table 9: Detail information about the 14 used datasets.
Dataset Task Image Number Label Type Range
LIVE (Sheikh et al., 2006) SDN-IQA 779 DMOS [1, 100]
CSIQ (Larson & Chandler, 2010) 866 DMOS [0, 1]
TID2013 (Ponomarenko et al., 2013) 3,000 MOS [0, 9]
KADID-10k (Lin et al., 2019) 10,125 MOS [1, 5]
SPAQ (Fang et al., 2020) ADN-IQA 11,125 MOS [0, 100]
LIVEC (Ghadiyaram & Bovik, 2015) 1,162 MOS [1, 100]
KonIQ-10K (Hosu et al., 2020) 10,073 MOS [0, 100]
GFIQA20k (Su et al., 2023a) F-IQA 19,988 MOS [0, 1]
AGIQA3k (Li et al., 2023a) AG-IQA 2,982 MOS [0, 1]
UWIQA (Yang et al., 2021a) U-IQA 890 MOS [0, 1]
AVA (Murray et al., 2012) IAA 250,000 MOS [0, 10]
AADB (Kong et al., 2016) IAA 10,000 MOS [0, 1]
exBeDDE (Zhao et al., 2020) D-IQA 1670 MOS [0, 1]
ECIQAD (Yue et al., 2023) EC-IQA 2400 MOS [1, 9]
Table 10: Detail information about the 14 used datasets.
Method Trainable Parms FLOPs Inference time KonIQ SRCC KADID SRCC
Q-Align (Wu et al., 2023) 8.2B (8200M) - 0.1s 0.938 0.934
LIQE (Zhang et al., 2023) 151M 17.40G 0.02s 0.919 0.930
Gamma 122.8M 28.45G 0.025s 0.939 0.962
A.4 Details of the Scene-based Differential Prompt
In the Scene-based Differential Prompt, we use different prompts for datasets from different scene. Specifically, we divide datasets into five categories, i.e., natural IQA, AI-generated IQA, underwater IQA, face IQA, natural IAA. We present the details in Table 11.
Table 11: Text prompts used in Scene-based Differential Prompt.
Dataset Prompt
LIVE, CSIQ, TID2013, KADID {natural bad-quality image, natural poor-quality image,
natural fair-quality image,
LIVEC, KonIQ, SPAQ natural good-quality image, natural perfect-quality image}
AGIQA3k {AI-generated bad-quality image, AI-generated poor-quality image,
AI-generated fair-quality image,
AI-generated good-quality image, AI-generated perfect-quality image}
GFIQA20k {face bad-quality image, face poor-quality image,
face fair-quality image,
face good-quality image, face perfect-quality image}
UWIQA {underwater bad-quality image, underwater poor-quality image,
underwater fair-quality image,
underwater good-quality image, underwater perfect-quality image}
AVA, AADB {natural bad-aesthetics image, natural poor-aesthetics image,
natural fair-aesthetics image,
natural good-aesthetics image, natural perfect-aesthetics image}
Appendix B Comparison with other mixed training methods
Table 12: Comparison with LIQE (Zhang et al., 2023) and UNIQUE (Zhang et al., 2021) when using the same training data.
Dataset LIVE CSIQ KADID BID LIVEC KonIQ Average
Method SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
UNIQUE 0.961 0.952 0.902 0.921 0.884 0.885 0.852 0.875 0.854 0.884 0.895 0.900 0.891 0.903
LIQE 0.970 0.951 0.936 0.939 0.930 0.931 0.875 0.900 0.904 0.910 0.919 0.908 0.922 0.923
Gamma 0.960 0.947 0.936 0.957 0.955 0.956 0.901 0.925 0.890 0.915 0.933 0.946 0.929 0.941
Table 13: Cross-dataset validation when using the same training data as LIQE and UNIQUE. The subscripts “$s$” and “$r$” stand for models trained on KADID and KonIQ, respectively. Gamma uses the same 6 datasets as LIQE and UNIQUE for training and validation. Gamma†† uses 12 datasets (include TID2012 and SPAQ) for training and performs zero-shot cross dataset validation on other datasets. Gamma†† significantly enhances the zero-shot performance on other datasets, showing the effectiveness of mix-training via MoAE and SDP methods.
Dataset TID2013 SPAQ AIGC2023 Average
NIQE 0.314 0.578 - 0.446
DBCNNs 0.686 0.412 0.730 0.609
PaQ2PiQ 0.423 0.823 0.643 0.630
MUSIQr 0.584 0.853 0.736 0.724
UNIQUE 0.768 0.838 0.761 0.789
LIQE 0.811 0.881 0.744 0.812
Gamma 0.805 0.894 0.770 0.823
Gamma†† - - 0.818 -
In this subsection, we conduct a more detailed comparison with other mixed training methods. We first compare with LIQE and UNIQUE using the same training data and data splitting ratios. As shown in Table 12, our method achieves better performance on most datasets than LIQE and UNIQUE, especially on the KADID (+2.5% SRCC) and BID (+2.6% SRCC) datasets compared with LIQE. On other datasets, i.e., LIVE and LIVEC, our model also achieves competitive results. Overall, our model has superior performance on these five datasets.
In addition, we conduct cross dataset validation under this setting. As shown in Table 13, our method achieves highly competitive results on TID2013 and SPAQ, demonstrating the strong generalization capability of our method. Note that using mix-training of more dataset can significantly enhance the zero-shot performance on other datasets, i.e., 0.770 SRCC to 0.818 SRCC on AIGC2023 dataset, demonstrating the effectiveness of mix-training via MoAE and SDP methods.
Compared with Q-Align, as shown in Table 14, our method achieves better results on KonIQ and KADID, and is also highly competitive on SPAQ.
Table 14: Comparison with Q-Align (Wu et al., 2023) when using the same training data.
Dataset KonIQ SPAQ KADID Average
Method SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
Q-Align 0.938 0.945 0.931 0.933 0.934 0.935 0.934 0.938
Gamma 0.940 0.950 0.928 0.932 0.962 0.964 0.943 0.949
Appendix C Sensitivity analysis of prompt
We analyze the sensitivity of prompts when the model is trained with scene-based differential prompts (SDP). Table 15 shows that using prompts different from SDP slightly reduces performance on most datasets, showing the robustness of our method. The quality prompt performs better than the general prompt on the IQA task, but performs worse on the IAA task, indicating the importance of appropriate prompts. In conclusion, our method is robust and insensitive to prompts, nevertheless we suggest using correct prompts to obtain better performance.
Table 15: Sensitivity analysis of prompt. Quality prompt is {bad-quality, poor-quality, fair-quality, good-quality, perfect-quality}; General prompt replaces the scene prompt (see Table 11) to “general”, e.g., {underwear bad-quality image} to {general bad-quality image}.
Dataset LIVEC KonIQ LIVE CSIQ AGIQA3k UWIQA AVA
Prompt SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
General prompt 0.882 0.888 0.921 0.920 0.943 0.930 0.948 0.957 0.775 0.843 0.832 0.842 0.648 0.624
Quality prompt 0.885 0.889 0.931 0.940 0.950 0.946 0.946 0.951 0.822 0.872 0.861 0.876 0.451 0.455
SDP 0.891 0.914 0.939 0.950 0.953 0.953 0.960 0.968 0.887 0.923 0.873 0.884 0.750 0.749
Appendix D Model efficiency analysis
We calculate the number of parameters, computation, and inference time of our model. For inference time, we use a $224\times 224$ image for testing. All indicators are obtained on a 3090 GPU. We compare it with two classic mixed training methods, LIQE (Zhang et al., 2023) and Q-Align (Wu et al., 2023). As shown in Table 10, our model achieves the best accuracy and efficiency. Compared with LIQE, our model has significantly better performance. Compared with Q-Align, we not only have better performance, but also have significantly lower model parameters and inference latency.

### v2

@@
 Gamma: Toward Generic Image Assessment with Mixture of Assessment Experts
-Abstract
+Abstract.
 Image assessment aims to evaluate the quality and aesthetics of images and has been applied across various scenarios, such as natural and AIGC scenes.
@@
 Extensive experiments show that our unified Gamma outperforms other state-of-the-art mixed-training methods by significant margins while covering more scenes.
-Machine Learning, ICML
-1 Introduction
+Codes are available at https://github.com/zht8506/Gamma.
+Generic Image Assessment, Unified Training, Mixture of Experts
+journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33nd ACM International Conference on Multimedia; October 27-31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33nd ACM International Conference on Multimedia (MM ’25), October 27-31, 2025, Dublin, Irelanddoi: XXXX.XXXXisbn: ccs: Computing methodologies Computer vision
+1. Introduction
 Image assessment is a long-standing research topic in the field of image processing, primarily comprising two tasks: Image Quality Assessment (IQA) and Image Aesthetic Assessment (IAA).
 These tasks require models to automatically evaluate the visual quality and aesthetic appeal of images, respectively.
-They have broad applications in various real-world scenarios, such as guiding image dehazing (Zhao et al., 2021), selecting high-quality images in a data engine (Rombach et al., 2022), serving as tools in an agentic system (Yang et al., 2024), or acting as reward models when aligning image generative models with human feedback (Liang et al., 2024).
+They have broad applications in various real-world scenarios, such as guiding image dehazing (Zhao et al., 2021), selecting high-quality images in a data engine (Rombach et al., 2022), serving as tools in an agentic system (Yang et al., 2024a), or acting as reward models when aligning image generative models with human feedback (Liang et al., 2024).
 Due to differences in image content and application scenarios, image assessment has spawned many sub-tasks, such as Natural-IQA for natural images, Face-IQA for facial images, AIGC-IQA for generated images, and IAA.
-Accordingly, numerous methods (Ke et al., 2021; He et al., 2022; Su et al., 2023b) have been proposed to address these specific tasks.
+Accordingly, numerous methods (Ke et al., 2021; He et al., 2022; Su et al., 2023a) have been proposed to address these specific tasks.
 However, these models often struggle to apply directly to other scenes or typically require task-specific fine-tuning on a given dataset.
 As illustrated in Figure 1, it is challenging for DEIQT (Qin et al., 2023) to transfer to other datasets without fine-tuning.
-This limitation prevents image assessment models from being widely applicable, e.g., IQA models are needed to assess facial, artistic, and natural images in the AIGC scene.
+This limitation prevents image assessment models from being widely applicable, e.g., IQA models are needed to assess facial, artistic, and natural images in the AIGC scene (Rombach et al., 2022; Yang et al., 2024a; Liang et al., 2024).
 Hence, there is an urgent need for a model that can effectively handle a variety of scenarios.
-Figure 1: “DEIQT Finetune” is trained and tested on the same dataset.
+Figure 1.
+“DEIQT Finetune” is trained and tested on the same dataset.
 “DEIQT ZS” directly assesses images using the model trained on other datasets, which perform poorly.
@@
 Our Gamma performs well on multiple datasets simultaneously and even surpass the task-specific method DETQT.
-Figure 2: Images with similar MOS labels from different datasets exhibit drastically different perceptual quality.
-It is not hard to observe that image (a) has clearly superior quality than the other three.
-Zoom in for a better view.
 To mitigate this issue, some approaches attempt to combine many datasets from various assessment tasks to train a general image assessment model.
-For instance, UNIQUE (Zhang et al., 2021) and LIQE (Zhang et al., 2023) utilize multiple authentic and synthetic natural IQA datasets for mixed training, but they focus only on Natural-IQA.
-Q-Align (Wu et al., 2023) uses a large language model with billions of parameters to unify IQA and IAA tasks, but it has a low inference speed and focuses solely on natural images.
-Additionally, PromptIQA (Chen et al., 2024b) employs image-score pairs as prompts for quality predictions, but it is inflexible as it requires multiple images as references during inference.
-These methods, however, fail to achieve competitive performance compared to models trained on specific datasets and cover a broad range of scenes.
-Thus, it is imperative to develop a foundational image assessment model capable of evaluating images from various scenes.
-To this end, we present Gamma, a Generic imAge assessMent model using Mixture of Assessment experts, to achieve unified image assessment across multiple datasets effectively.
+For instance, UNIQUE (Zhang et al., 2021) and LIQE (Zhang et al., 2023) utilize multiple authentic and synthetic natural IQA datasets for mixed training, but they focus only on natural image.
+Q-Align (Wu et al., 2023) uses a large language model with billions of parameters to unify IQA and IAA tasks, but it has a low inference speed and still solely focuses on natural images.
+PromptIQA (Chen et al., 2024b) employs image-score pairs as prompts for quality predictions, but it is inflexible as it requires multiple images as references during inference.
+These methods fail to cover a broad range of scenes and achieve competitive performance compared to models trained on specific datasets.
+Thus, it is imperative to develop an effective and generic image assessment model capable of evaluating images from various scenes.
+To this end, we present Gamma, a Generic imAge assessMent model using Mixture of Assessment experts, to achieve unified image assessment across multiple scenes effectively.
 We found that the primary challenge in mixed-dataset training is the mean opinion score (MOS) bias between different datasets, e.g., images with similar MOS may exhibit different perceptual qualities across various datasets.
@@
 To address this challenge, we introduce a novel Mixture of Assessment Experts (MoAE) module in our Gamma model.
-MoAE consists of two types of experts: shared experts and adaptive experts.
+It consists of two types of experts: shared experts and adaptive experts.
 The shared experts are employed throughout to learn dataset-shared knowledge, while the adaptive experts are dynamically activated to varying degrees to learn dataset-specific knowledge.
 Additionally, an image-based router modulates the contributions of each adaptive expert.
-This strategy allows the model to capture common features while flexibly adjusting representative features for different datasets.
-Compared with general Mixture of Experts (MoE) (Shazeer et al., 2017) and Lora-based MoE (Liu et al., 2024), we equip the MoAE module only in the rear blocks instead of all blocks, making it more efficient.
+The MoAE allows the model to capture common features while flexibly adjusting representative features for different datasets.
+Compared with general Mixture of Experts (MoE) (Shazeer et al., 2017) and LoRA-based MoE (Liu et al., 2024b), we equip the MoAE only in the rear blocks instead of all blocks, making it more efficient.
 Secondly, we propose a Scene-based Differential Prompt (SDP), which uses different prompts for different datasets according to their scenes.
 This strategy provides scene-specific knowledge for representation learning of different datasets, guiding the mixed-dataset training process.
+Figure 2.
+Images with similar MOS labels from different datasets exhibit drastically different perceptual quality.
+It is not hard to observe that image (a) has clearly superior quality than the other three.
+Zoom in for a better view.
 Our Gamma model is uniformly trained on a mixture of 12 datasets from 6 image assessment scenarios spanning IQA and IAA tasks.
@@
   • Extensive experiments show that Gamma achieves SOTA performance on 12 datasets across 6 image assessment scenes in both mixed training and task-specific training settings.
-2 Related Work
-2.1 Image Assessment
+2. Related Work
+2.1. Image Assessment
 Image Assessment mainly includes two subtasks: Image Quality Assessment (IQA) and Image Aesthetic Assessment (IAA).
 In the deep learning era, these two tasks have achieved significant breakthroughs.
-For the IQA task, researchers develop various advanced techniques to improve performance, including multi-level feature aggregation (Li et al., 2018; Chen et al., 2024a; Xu et al., 2024; Zhang et al., 2018; Mittal et al., 2012b; Ying et al., 2020), adaptive convolution (Su et al., 2020), transformer methods (Ke et al., 2021; Qin et al., 2023), vision-language models (VLMs) (Wang et al., 2023; Zhang et al., 2023) and large language models (LLM) (You et al., 2023).
-Moreover, besides the natural image assessment, these are various IQA methods for other scenes, such as face IQA (Ou et al., 2021; Su et al., 2023b; Jo et al., 2023), AIGC IQA (Yuan et al., 2023), underwater IQA (Yang et al., 2021b; Guo et al., 2023; Yang & Sowmya, 2015; Liu et al., 2023).
-For the IAA task, numerous methods have also been proposed to improve performance, including loss function (Talebi & Milanfar, 2018), novel transformer architecture (Tu et al., 2022), multi-level features (Hosu et al., 2019), theme information (He et al., 2022; Li et al., 2023b) and multimodal pre-training (Ke et al., 2023).
-2.2 Mixed Training for Image Assessment
+For the IQA task, researchers develop various advanced techniques to improve performance, including feature and model structure design (Ying et al., 2020; Zhang et al., 2018), multi-level feature aggregation (Li et al., 2018; Chen et al., 2024a; Xu et al., 2024), adaptive convolution (Su et al., 2020), transformer methods (Ke et al., 2021; Qin et al., 2023), vision language models (VLMs) (Wang et al., 2023; Zhang et al., 2023) and large language models (LLM) (You et al., 2023; Wu et al., 2023).
+Moreover, besides the natural image assessment, these are various IQA methods for other scenes, such as face IQA (Ou et al., 2021; Su et al., 2023a; Jo et al., 2023), AIGC IQA (Yuan et al., 2023), underwater IQA (Yang et al., 2021b; Guo et al., 2023; Yang and Sowmya, 2015; Liu et al., 2023).
+For the IAA task, numerous methods have also been proposed to improve performance, including loss function (Talebi and Milanfar, 2018), novel transformer architecture (Tu et al., 2022), multi-level features (Hosu et al., 2019), theme information (He et al., 2022; Li et al., 2023a) and multimodal pre-training (Ke et al., 2023).
+Figure 3.
+(a) The architecture of Gamma: It consists of a visual encoder $\mathcal{V}$ and text encoder $\mathcal{T}$; We add the Mixture of Assessment Experts (MoAE) to the last few layers of both encoders.
+We introduce a Scene-based Differential Prompt (SDP) to prompt images from different scenes (See Section 3.4 for details).
+(b) The MoAE module: It involves one shared expert ($E^{\mathrm{shared}}$) and several adaptive experts (from $E_{1}^{\mathrm{adaptive}}$ to $E_{n}^{\mathrm{adaptive}}$).
+We employ a router to adaptively activate the adaptive experts.
+We then use a learnable factor $\sigma$ to merge the features of two type of experts.
+2.2. Mixed Training for Image Assessment
 As a fundamental image processing task, image assessment has been effectively applied to various scenarios.
@@
 Our Gamma outperforms other SOTA mixed-training methods by significant margins while covering more scenes.
-2.3 Mixture of Experts
+2.3. Mixture of Experts
 Mixture-of-Experts (MoE) divides specific parts of the parameters into several subsets, each of which is called an expert.
@@
 In Gamma, we utilize novel MoE module to effectively learn specific and general features of multi-dataset.
-Figure 3: (a) The architecture of Gamma: It consists of a visual encoder $\mathcal{V}$ and text encoder $\mathcal{T}$; We add the Mixture of Assessment Experts (MoAE) to the last few layers of both encoders.
-We introduce a Scene-based Differential Prompt (SDP) to prompt images from different scenes (See Section 3.4 for details).
-(b) The MoAE module: It involves one shared expert ($E^{\mathrm{shared}}$) and several adaptive experts (from $E_{1}^{\mathrm{adaptive}}$ to $E_{n}^{\mathrm{adaptive}}$).
-We employ a router to adaptively activate the adaptive experts.
-We then use a learnable factor $\sigma$ to merge the features of two type of experts.
-3 Method
-3.1 Preliminary
-As a foundational vision-language model (VLM), CLIP (Radford et al., 2021) has shown significant promise in supporting a wide array of vision tasks.
-Specifically, CLIP is composed of a transformer-based visual encoder $\mathcal{V}$ and a text encoder $\mathcal{T}$, which generate aligned visual representations $I$ and text representations $T$ for each image-text pair.
+3. Method
+3.1. Preliminary
+As a foundational vision-language model (VLM), CLIP (Radford et al., 2021) has shown significant promise in supporting a wide array of vision tasks, including IQA and IAA tasks (Wang et al., 2023; Zhang et al., 2024; Pan et al., 2024).
+Specifically, CLIP is composed of a visual encoder $\mathcal{V}$ and a text encoder $\mathcal{T}$, which generate aligned visual representations $I$ and text representations $T$ for each image-text pair.
 Utilizing these features, we can compute cosine similarity scores between image and text pairs across different domains or tasks to perform task-specific predictions, including image assessment tasks.
@@
 In the following, we will provide a detailed exposition of its components.
-3.2 Overview of Gamma
+Figure 4.
+Visual examples from different datasets, which include natural images, underwater images, face images, AIGC images, and etc.
+The first value is the prediction score and the second value is the ground-truth MOS.
+Our Gamma accurately evaluates images from different scenes, demonstrating the generalization and effectiveness.
+All images are resized for better visibility.
+3.2. Overview of Gamma
 As illustrated in Figure 3, our Gamma employs a visual encoder $\mathcal{V}$ to extract visual features $I\in\mathbb{R}^{d}$, and a text encoder $\mathcal{T}$ to extract text features $T\in\mathbb{R}^{d}$.
@@
 This process can be described as:
-$\mathrm{q}=\sum_{k=1}^{5}\mathrm{C}_{k}\mathrm{Softmax}({I^{\prime}}^{\top}T_{k}/\tau),I^{\prime}=\mathrm{Adapter}(I),$ (1)
+(1) $\mathrm{q}=\sum_{k=1}^{5}\mathrm{C}_{k}\mathrm{Softmax}({I^{\prime}}^{\top}T_{k}/\tau),I^{\prime}=\mathrm{Adapter}(I),$
 […]
-3.3 Mixture of Assessment Experts
+3.3. Mixture of Assessment Experts
 To develop a unified and generic image assessment model, we aim to combine multiple image assessment datasets for joint training.
@@
 To address this challenge, we propose the MoAE module, where several experts are employed to learn the diverse biases of different datasets.
-As shown in Figure 3 (b), the proposed MoAE module includes a shared assessment expert ($E^{\mathrm{shared}}$) to learn common knowledge of image assessment and several adaptive assessment experts (from $E_{1}^{\mathrm{adaptive}}$ to $E_{n}^{\mathrm{adaptive}}$) to dynamically learn dataset-specific knowledge.
+As shown in Figure 3 (b), the MoAE includes a shared assessment expert ($E^{\mathrm{shared}}$) to learn common knowledge of image assessment and several adaptive assessment experts (from $E_{1}^{\mathrm{adaptive}}$ to $E_{n}^{\mathrm{adaptive}}$) to dynamically learn dataset-specific knowledge.
 The Shared Assessment Expert.
-The shared assessment expert $E^{\mathrm{shared}}$ inherits the image assessment capabilities of the original CLIP model by reusing its weights.
+The shared assessment expert $E^{\mathrm{shared}}$ inherits the image assessment capabilities of the original multimodal model (e.g., CLIP (Radford et al., 2021) or UniQA (Zhou et al., 2024b)) by reusing its weights.
 This expert remains frozen during training to ensure that the learned world knowledge is retained.
@@
 Given an input hidden state $x\in\mathbb{R}^{d}$, the output of the shared assessment expert is:
-$y^{\mathrm{shared}}=E^{\mathrm{shared}}(x),$ (2)
+(2) $y^{\mathrm{shared}}=E^{\mathrm{shared}}(x),$
 where $E^{\mathrm{shared}}(\cdot)$ is implemented as the original feed-forward network (FFN) of the CLIP model.
-Figure 4: Visual examples from different datasets, which include natural images, underwater images, face images, AI-generated images, and etc.
-The first value is the prediction score and the second value is the ground-truth MOS.
-Our Gamma can accurately evaluate images from different scenes, demonstrating the generalization and effectiveness.
-All images are resized for better visibility.
 The Adaptive Assessment Expert.
@@
 Given an input feature $x\in\mathbb{R}^{d}$, the output $y^{\mathrm{adaptive}}$ can be computed as:
-$y^{\mathrm{adaptive}}$ $=\sum_{i=1}^{n}G(x)_{i}E_{i}^{\mathrm{adaptive}}(x),$ (3)
+(3) $y^{\mathrm{adaptive}}$ $=\sum_{i=1}^{n}G(x)_{i}E_{i}^{\mathrm{adaptive}}(x),$
 $G(x)$ $=\mathrm{Softmax}(Wx).$
 Here, the router $G(\cdot)$ is a linear transformation for the input feature $x$; $W\in\mathbb{R}^{n\times d}$ is the transformation matrix.
-To avoid unreasonable weights, we utilize a $\mathrm{Softmax}$ operator to normalize the contribution weights.
-This ensures that the model can learn dataset-specific knowledge efficiently.
-Notably, we initialize these experts with the weights of pre-trained CLIP to equip them with image assessment capabilities.
+We utilize a $\mathrm{Softmax}$ operator to normalize the contribution weights.
+Notably, we initialize these experts with the weights of pre-trained multimodal model to equip them with preliminary image assessment capabilities.
 MoAE Module.
@@
 Thus, the output of the MoAE module can be expressed as:
-$y^{\mathrm{MoAE}}=y^{\mathrm{shared}}+\sigma\cdot y^{\mathrm{adaptive}}.\$ (4)
-The $\sigma$ factor is zero-initialized so that the visual and text encoders can generate aligned features at the beginning.
-In practice, we freeze the shared experts and set the adaptive experts to be tunable only.
-This approach maintains parameter efficiency during mixed training and preserves the multi-modal capabilities of the original model.
-We incorporate the MoAE module into the last $K$ layers of both visual and text encoders, as shown in Figure 3.
+(4) $y^{\mathrm{MoAE}}=y^{\mathrm{shared}}+\sigma\cdot y^{\mathrm{adaptive}}.\$
+The $\sigma$ factor is zero-initialized so that the visual and text encoders can generate aligned features at the beginning of training.
+We freeze the shared experts and set the adaptive experts to be tunable only.
+This approach preserves the learned knowledge of the original model and improves training efficiency.
+Inspired by (Yang et al., 2024b), which points out that higher layers are easier to tune for downstream tasks and freezing lower layers preserves more generalizable knowledge, we incorporate the MoAE into the last $K$ layers of both visual and text encoders, as shown in Figure 3.
 This strategy makes our method both effective and efficient.
 The visual feature extraction process can be formulated as:
-$I_{i}$ $=\bm{\mathcal{V}}_{i}(I_{i-1}),\quad i=1,2,\ldots,L-K$ (5)
+(5) $I_{i}$ $=\bm{\mathcal{V}}_{i}(I_{i-1}),\quad i=1,2,\ldots,L-K$
 $I_{j}$ $=\bm{\mathcal{V}}_{j}^{\mathrm{MoAE}}(I_{j-1}),\quad j=L-K+1,\ldots,L$
@@
 The text branch operates similarly to the visual branch.
-3.4 Scene-based differential prompt
-To facilitate scene-guided learning, we introduce a Scene-based Differential Prompt (SDP) to help the model acquire scene-specific knowledge from different datasets.
-We utilize 12 datasets spanning 6 image assessment scenarios, including synthetic distortion nature IQA, authentic distortion nature IQA, face IQA, AIGC IQA, underwater IQA, and IAA, for mixed training.
-We categorize these datasets into five groups based on their scenes: natural quality, AI-generated quality, underwater quality, face quality, and natural aesthetics.
-Specifically, for the face quality assessment dataset, we use prompts such as face bad-quality, face poor-quality, face fair-quality, face good-quality, face perfect-quality appended to the word image.
-For more details on these prompts, please refer to Appendix A.4.
-This strategy effectively differentiates the feature space of images from various scenes and enhances scene-specific knowledge, thereby mitigating the MOS bias across different datasets.
-Experimental results show that this method significantly improves the model’s performance (Table 1).
-4 Experiments
-4.1 Datasets
-We utilize 12 datasets for unified training and testing, encompassing both image quality and aesthetic assessment tasks.
-For the IQA task, five different assessment scenarios are included: synthetic distortion nature IQA (SDN-IQA), authentic distortion nature IQA (ADN-IQA), face IQA (F-IQA), AIGC IQA (AG-IQA), and underwater IQA (U-IQA).
-For the IAA task, we use two classical benchmarks, AVA and AADB.
-In addition, we use two rare datasets to verify the generalization ability of the model, i.e., exBeDDE and ECIQAD.
-The exBeDDE is a dehazed image quality assessment (D-IQA) dataset, and the ECIQAD is an enhanced colonoscopy image quality assessment (EC-IQA) dataset.
-Details about these datasets is provided in Table 9.
-Table 1: Comparison with SOTA task-specific and mixed-training models on 12 datasets for 6 image assessment tasks.
+Table 1.
+Comparison with SOTA task-specific and mixed-training models on 12 datasets for 6 image assessment tasks.
 “Gamma” and “Gamma-T” denote the mixed-training and task-specific models, respectively.
@@
 indicates that we retrain the model with the same data split as ours.
-Task Synthetic Distortion Nature IQA (SDN-IQA) Authentic Distortion Nature IQA (ADN-IQA)
+Task Synthetic Distortion Natural IQA (SDN-IQA) Authentic Distortion Natural IQA (ADN-IQA)
 Dataset LIVE CSIQ TID2013 KADID LIVEC KonIQ SPAQ
@@
 Training Type Method SRCC PLCC Method SRCC PLCC Method SRCC PLCC Method SRCC PLCC Method SRCC PLCC
-SDD-FIQA 0.602 0.649 DBCNN 0.821 0.876 FDUM 0.694 0.689 MaxViT 0.708 0.745 MUSIQ 0.706 0.712
-Task Specific IFQA 0.697 0.722 HyperNet 0.836 0.890 UCIQE 0.627 0.626 TANet 0.758 0.765 TANet 0.738 0.737
+IFQA 0.960 0.960 DBCNN 0.821 0.876 FDUM 0.694 0.689 MaxViT 0.708 0.745 MUSIQ 0.706 0.712
+Task Specific StyleGAN 0.967 0.968 HyperNet 0.836 0.890 UCIQE 0.627 0.626 TANet 0.758 0.765 TANet 0.738 0.737
 TOPIQ 0.966 0.967 CLIPIQA 0.843 0.805 URanker 0.674 0.663 VILA 0.774 0.774 TAVAR 0.761 0.763
-GPFIQA 0.964 0.965 PSCR 0.850 0.906 UIQI 0.742 0.741 UniQA 0.776 0.776 UniQA 0.786 0.787
+CR-FIQA 0.959 0.960 MUSIQ 0.826 0.866 UIQI 0.742 0.741 ELTA 0.764 0.777 ELTA 0.760 0.772
+GPFIQA 0.964 0.965 PSCR 0.850 0.906 HPUIQA 0.796 0.790 UniQA 0.776 0.776 UniQA 0.786 0.787
 Gamma-T 0.968 0.968 Ours-T 0.894 0.921 Ours-T 0.870 0.880 Ours-T 0.785 0.784 Ours-T 0.793 0.798
@@
 Gamma 0.970 0.970 Gamma 0.887 0.923 Gamma 0.873 0.884 Gamma 0.750 0.749 Gamma 0.756 0.755
-Table 2: The effect of Mixture of Assessment Experts (MoAE) and Scene-based Differential Prompt (SDP).
-Dataset LIVEC KonIQ LIVE CSIQ AGIQA3k UWIQA AVA
-MoAE SDP SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
-$\times$ $\times$ 0.765 0.792 0.858 0.885 0.927 0.918 0.852 0.898 0.800 0.866 0.750 0.768 0.681 0.672
-$\times$ $\checkmark$ 0.843 0.856 0.874 0.896 0.929 0.917 0.866 0.901 0.841 0.887 0.770 0.780 0.721 0.715
-$\checkmark$ $\times$ 0.851 0.871 0.940 0.949 0.957 0.952 0.949 0.966 0.870 0.910 0.863 0.878 0.740 0.737
-$\checkmark$ $\checkmark$ 0.891 0.914 0.939 0.950 0.960 0.968 0.953 0.953 0.887 0.923 0.873 0.884 0.750 0.749
-Table 3: The impact of different number of experts in adaptive experts. We use naive prompt strategy for ablation.
-Dataset LIVEC CSIQ TID2013 AGIQA3k UWIQA
-Experts Number SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
-Zero Expert 0.765 0.792 0.852 0.898 0.792 0.826 0.800 0.866 0.750 0.768
-One Expert 0.842 0.866 0.945 0.963 0.918 0.931 0.866 0.908 0.859 0.873
-Three Experts 0.851 0.871 0.949 0.966 0.926 0.934 0.870 0.910 0.863 0.878
-Five Experts 0.854 0.889 0.951 0.965 0.926 0.934 0.872 0.911 0.860 0.876
-4.2 Implementation Details
+3.4. Scene-based differential prompt
+To facilitate scene-guided learning, we introduce a Scene-based Differential Prompt (SDP) to help the model acquire scene-specific knowledge from different datasets.
+We utilize 12 datasets spanning 6 image assessment scenarios, including synthetic distortion Natural IQA, authentic distortion Natural IQA, face IQA, AIGC IQA, underwater IQA, and IAA, for mixed training.
+We categorize these datasets into five groups based on their scenes: natural quality, AI-generated quality, underwater quality, face quality, and natural aesthetics.
+Specifically, for the face quality assessment dataset, we use prompts such as face bad-quality, face poor-quality, face fair-quality, face good-quality, face perfect-quality appended to the word image.
+Then, the text prompt is sent to the text encoder of the Gamma to obtain text features.
+Text features will be combined with image features for quality prediction (Section 3.2).
+For more details on these prompts, please refer to Supplementary Materials.
+This strategy effectively differentiates the feature space of images from various scenes and enhances scene-specific knowledge, thereby mitigating the MOS bias across different datasets.
+Experimental results show that this method significantly improves the model’s performance (Table 1).
+4. Experiments
+4.1. Datasets
+We utilize 12 datasets for unified training and testing, encompassing both IQA and IAA tasks.
+For the IQA task, five different assessment scenarios are included: synthetic distortion natural IQA (SDN-IQA), authentic distortion natural IQA (ADN-IQA), face IQA (F-IQA), AIGC IQA (AG-IQA), and underwater IQA (U-IQA).
+For the IAA task, we use two classical benchmarks, AVA and AADB.
+In addition, we use two rare datasets to verify the generalization ability of the model, i.e., exBeDDE and ECIQAD.
+The exBeDDE is a dehazed image quality assessment (D-IQA) dataset, and the ECIQAD is an enhanced colonoscopy image quality assessment (EC-IQA) dataset.
+Details about these datasets is provided in supplementary materials.
+4.2. Implementation Details
 Following the settings in (Su et al., 2020; Ke et al., 2021), we randomly divide each dataset into 80% for training and 20% for testing.
@@
 For the task-specific training, we use different training settings according to the task and size of datasets.
-More training details are provided in the appendix.
-Table 4: The impact of different model configuration in the proposed MoAE module.
+More training details are provided in the supplementary materials.
+Table 2. The effect of Mixture of Assessment Experts (MoAE) and Scene-based Differential Prompt (SDP).
+Dataset LIVEC KonIQ LIVE CSIQ AGIQA3k UWIQA AVA
+MoAE SDP SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
+$\times$ $\times$ 0.765 0.792 0.858 0.885 0.927 0.918 0.852 0.898 0.800 0.866 0.750 0.768 0.681 0.672
+$\times$ $\checkmark$ 0.843 0.856 0.874 0.896 0.929 0.917 0.866 0.901 0.841 0.887 0.770 0.780 0.721 0.715
+$\checkmark$ $\times$ 0.851 0.871 0.940 0.949 0.957 0.952 0.949 0.966 0.870 0.910 0.863 0.878 0.740 0.737
+$\checkmark$ $\checkmark$ 0.891 0.914 0.939 0.950 0.960 0.968 0.953 0.953 0.887 0.923 0.873 0.884 0.750 0.749
+Table 3. The impact of different number of experts in adaptive experts. We use naive prompt strategy for ablation.
+Dataset LIVEC CSIQ TID2013 AGIQA3k UWIQA Average
+Experts Number SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
+Zero Expert 0.765 0.792 0.852 0.898 0.792 0.826 0.800 0.866 0.750 0.768 0.792 0.83
+One Expert 0.842 0.866 0.945 0.963 0.918 0.931 0.866 0.908 0.859 0.873 0.886 0.908
+Three Experts 0.851 0.871 0.949 0.966 0.926 0.934 0.870 0.910 0.863 0.878 0.892 0.912
+Five Experts 0.854 0.889 0.951 0.965 0.926 0.934 0.872 0.911 0.860 0.876 0.893 0.915
+Table 4. The impact of different model configuration in the proposed MoAE module.
 Dataset LIVEC CSIQ AGIQA3k UWIQA AVA
@@
 Our MoAE module 0.851 0.871 0.949 0.966 0.870 0.910 0.863 0.878 0.740 0.737
-Table 5: The impact of adding MoAE to different numbers of layers for training. Time means training time.
-Dataset Parms Time LIVEC KonIQ LIVE CSIQ AGIQA3k UWIQA AVA
-MoE Layer Million Hours SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
-w/o MoAE 149.9 3.5 0.765 0.792 0.858 0.885 0.927 0.918 0.852 0.898 0.800 0.866 0.750 0.768 0.681 0.672
-Last 4 layers 231.8 7.5 0.830 0.859 0.933 0.944 0.954 0.952 0.937 0.960 0.866 0.909 0.853 0.867 0.735 0.732
-Last 6 layers 272.7 10.2 0.851 0.871 0.940 0.949 0.957 0.952 0.949 0.966 0.870 0.910 0.863 0.878 0.740 0.737
-Last 8 layers 313.6 13.4 0.852 0.883 0.941 0.947 0.956 0.951 0.953 0.967 0.872 0.913 0.866 0.875 0.746 0.743
-All 12 layers 395.5 17.2 0.860 0.883 0.939 0.950 0.954 0.950 0.954 0.968 0.881 0.908 0.863 0.868 0.728 0.725
-4.3 Main Results
+4.3. Main Results
 Our MoAE-equipped model can be used for task-specific training and mixed training, both of which can achieve state-of-the-art (SOTA) performance, as shown in Table 1.
+We conduct a comprehensive performance comparison of the Gamma with 30 SOTA IQA methods on the 6 IQA tasks.
+We compare natural scene IQA methods, including HyperIQA (Su et al., 2020), MUSIQ (Ke et al., 2021), TOPIQ (Chen et al., 2024a), DEIQT (Qin et al., 2023), LoDA (Xu et al., 2024), UniQA (Zhou et al., 2024b), UNIQUE (Zhang et al., 2021), LIQE (Zhang et al., 2023), StairIQA (Sun et al., 2023) and PromptIQA (Chen et al., 2024b); face IQA methods, including IFQA (Jo et al., 2023), StyleGAN-IQA from (Su et al., 2023a), CR-FIQA (Boutros et al., 2023), GPFIQA (Su et al., 2023a); AIGC IQA methods, including DBCNN (Zhang et al., 2018), CLIPIQA (Wang et al., 2023) and PSCR (Yuan et al., 2023); underwater IQA methods, including FDUM (Yang et al., 2021a), UIQI (Liu et al., 2023), URanker (Guo et al., 2023), UCIQE (Yang and Sowmya, 2015) and HPUIQA (Liu et al., 2023), and IAA methods, including VILA (Ke et al., 2023), TANet (He et al., 2022), MaxViT (Tu et al., 2022), TAVAR (Li et al., 2023a) and ELTA (Liu et al., 2024a).
 Task-specific Training.
@@
 We conduct mixed training on 12 image assessment datasets.
-The trained model can be used to assess the images from these datasets.
+The trained model can be used to assess the images from these datasets directly.
 The experimental results are reported in Table 1.
-When compared with other mixed training models, such as StairIQA and PromptIQA, our method exhibits powerful and superior performance on each dataset.
+When compared with other mixed training models, our method exhibits competitive results on each dataset.
 More importantly, our method can also be used to IAA tasks and demonstrates excellent performance.
@@
 These results demonstrate that our approach can be effectively applied to different image assessment scenarios.
-Comparison with other mixed training methods.
-We conduct a detailed comparison with other mixed training methods, including LIQE, UNIQUE and Q-Align, with the same training datasets.
-We also perform the cross-dataset validation with LIQE and UNIQUE.
-Please refer to Appendix B for datails.
 Qualitative Results.
@@
 These results shows the high generalization capability of our Gamma.
-4.4 Ablation Studies
-We conduct detailed ablation studies to validate the effectiveness of our proposed modules.
+Table 5. The impact of adding MoAE to different numbers of layers for training. Time means training time.
+Dataset Parms Time LIVEC KonIQ LIVE CSIQ AGIQA3k UWIQA AVA
+MoE Layer Million Hours SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
+w/o MoAE 149.9 3.5 0.765 0.792 0.858 0.885 0.927 0.918 0.852 0.898 0.800 0.866 0.750 0.768 0.681 0.672
+Last 4 layers 231.8 7.5 0.830 0.859 0.933 0.944 0.954 0.952 0.937 0.960 0.866 0.909 0.853 0.867 0.735 0.732
+Last 6 layers 272.7 10.2 0.851 0.871 0.940 0.949 0.957 0.952 0.949 0.966 0.870 0.910 0.863 0.878 0.740 0.737
+Last 8 layers 313.6 13.4 0.852 0.883 0.941 0.947 0.956 0.951 0.953 0.967 0.872 0.913 0.866 0.875 0.746 0.743
+All 12 layers 395.5 17.2 0.860 0.883 0.939 0.950 0.954 0.950 0.954 0.968 0.881 0.908 0.863 0.868 0.728 0.725
+4.4. Ablation Studies
+We conduct detailed ablation studies to validate the effectiveness of our modules.
 Note that we use naive prompt strategy (described in Section 3.2) to perform all ablations unless otherwise specified.
@@
 Considering the page limit, we only show the datasets with relatively large differences in results.
-More ablations about prompt sensitivity and model efficiency please refer to Appendix C and D.
+More ablations about prompt sensitivity and model efficiency please refer to Supplementary Materials.
+Table 6. Results when only one adaptive expert is activated. The weights factors of other experts are set to 0. It can be observed that different experts focus on different datasets.
+Dataset LIVEC KonIQ LIVE CSIQ AGIQA3k UWIQA GFIQA AVA
+Expert index SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
+$1$-th expert 0.847 0.860 0.927 0.938 0.933 0.933 0.894 0.906 0.815 0.870 0.770 0.779 0.959 0.957 0.666 0.673
+$2$-th expert 0.715 0.672 0.681 0.717 0.900 0.861 0.815 0.846 0.832 0.885 0.755 0.756 0.826 0.797 0.663 0.652
+$3$-th expert 0.768 0.741 0.794 0.818 0.918 0.917 0.833 0.877 0.808 0.910 0.691 0.709 0.903 0.897 0.715 0.716
+Gamma 0.851 0.871 0.940 0.949 0.957 0.952 0.949 0.966 0.870 0.910 0.863 0.878 0.970 0.970 0.740 0.737
+Table 7. Comparison with LIQE (Zhang et al., 2023) and UNIQUE (Zhang et al., 2021) when using the same training data.
+Dataset LIVE CSIQ KADID BID LIVEC KonIQ Average
+Method SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
+UNIQUE 0.961 0.952 0.902 0.921 0.884 0.885 0.852 0.875 0.854 0.884 0.895 0.900 0.891 0.903
+LIQE 0.970 0.951 0.936 0.939 0.930 0.931 0.875 0.900 0.904 0.910 0.919 0.908 0.922 0.923
+Gamma 0.960 0.947 0.936 0.957 0.955 0.956 0.901 0.925 0.890 0.915 0.933 0.946 0.929 0.941
+Table 8. Comparison with Q-Align (Wu et al., 2023) when using the same training data.
+Dataset KonIQ SPAQ KADID Average
+Method SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
+Q-Align 0.938 0.945 0.931 0.933 0.934 0.935 0.934 0.938
+Gamma 0.940 0.950 0.928 0.932 0.962 0.964 0.943 0.949
+Table 9.
+Generalization capability validation on the exBeDDE and ECIQAD datasets.
+The “Pretrained weight” denotes the model weight of mixed training.
+Loading pretrained weight for initialization improves model performance.
+(a) Results on the exBeDDE dataset.
+Method SRCC PLCC
+BRISQUE (Mittal et al., 2012) 0.890 0.906
+PSQA-I (Liu et al., 2019) 0.907 0.924
+HyperIQA (Su et al., 2020) 0.917 0.926
+FADE (Choi et al., 2015) 0.714 0.729
+DHQI (Min et al., 2018) 0.919 0.939
+VDA-DQA (Guan et al., 2022) 0.923 0.942
+Ours 0.916 0.938
+Ours + Pretrained weight 0.937 0.951
+(b) Results on the ECIQAD dataset.
+Method SRCC PLCC
+BRISQUE (Mittal et al., 2012) 0.436 0.459
+BIQME (Gu et al., 2017) 0.770 0.768
+BPRI (Min et al., 2017) 0.152 0.181
+FRIQUEE (Ghadiyaram and Bovik, 2017) 0.663 0.656
+CIQA (Chen et al., 2021) 0.738 0.735
+ECIQ (Ke et al., 2021) 0.839 0.842
+Ours 0.912 0.922
+Ours + Pretrained weight 0.917 0.927
 Effectiveness of the prompt strategy.
@@
 Therefore, the two methods are mutually beneficial.
-Figure 5: (a) Average performance when using different numbers of experts.
+Figure 5.
+(a) Average performance when using different numbers of experts.
 Adding adaptive experts significantly improves model performance.
-(b) Average activations of three experts in the last layer of the visual encoder with naive prompts.
+(b) Average activations of three experts in the last layer of the visual encoder.
 Image evaluations of different scenes have different activation patterns.
@@
 We use three experts to constitute the adaptive assessment experts in MoAE to achieve the optimal trade-off between accuracy and efficiency.
-Table 6: Results when only one adaptive expert is activated. The weights factors of other experts are set to 0. It can be observed that different experts focus on different datasets.
-Dataset LIVEC KonIQ LIVE CSIQ AGIQA3k UWIQA GFIQA AVA
-Expert index SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
-$1$-th expert 0.847 0.860 0.927 0.938 0.933 0.933 0.894 0.906 0.815 0.870 0.770 0.779 0.959 0.957 0.666 0.673
-$2$-th expert 0.715 0.672 0.681 0.717 0.900 0.861 0.815 0.846 0.832 0.885 0.755 0.756 0.826 0.797 0.663 0.652
-$3$-th expert 0.768 0.741 0.794 0.818 0.918 0.917 0.833 0.877 0.808 0.910 0.691 0.709 0.903 0.897 0.715 0.716
-Gamma 0.851 0.871 0.940 0.949 0.957 0.952 0.949 0.966 0.870 0.910 0.863 0.878 0.970 0.970 0.740 0.737
-Table 7: Generalization capability validation on the exBeDDE and ECIQAD datasets.
-The “Pretrained weight” denotes the model weight of mixed training.
-We can notice that loading pretrained weight for initialization can improve model performance.
-(a) Results on the exBeDDE datasets.
-Method SRCC PLCC
-BRISQUE (Mittal et al., 2012a) 0.890 0.906
-PSQA-I (Liu et al., 2019) 0.907 0.924
-HyperIQA (Su et al., 2020) 0.917 0.926
-FADE (Choi et al., 2015) 0.714 0.729
-DHQI (Min et al., 2018) 0.919 0.939
-VDA-DQA (Guan et al., 2022) 0.923 0.942
-Ours 0.916 0.938
-Ours + Pretrained weight 0.937 0.951
-(b) Results on the ECIQAD datasets.
-Method SRCC PLCC
-BRISQUE (Mittal et al., 2012a) 0.436 0.459
-BIQME (Gu et al., 2017) 0.770 0.768
-BPRI (Min et al., 2017) 0.152 0.181
-FRIQUEE (Ghadiyaram & Bovik, 2017) 0.663 0.656
-CIQA (Chen et al., 2021) 0.738 0.735
-ECIQ (Ke et al., 2021) 0.839 0.842
-Ours 0.912 0.922
-Ours + Pretrained weight 0.917 0.927
 Effect of freezing the shared expert.
@@
 These results show that it is beneficial to utilize aligned features at the beginning of training and partially using features from adaptive experts.
+We plot the value change of $\sigma$ during training in Figure 6.
+The absolute value of $\sigma$ increases gradually and eventually stabilize.
+This means that the adaptive expert is constantly fitting the bias between different datasets.
+The value is negative, suggesting that the adaptive experts remove the preference scores from the shared expert, similar to a denoising process.
 Adding adapter to last few layers.
@@
 We visualize the average activation degree of three experts in the last layer of Gamma’s visual encoder for different datasets, as shown in Figure LABEL:fig:ab-activation.
-We can observe that the activation patterns are different for different scenarios.
-Specifically, the natural image assessment datasets, e.g., LIVE, CSIQ, KADID, show different activation patterns from the face IQA dataset GFIQA and the underwater IQA dataset UWIQA.
-The synthetic distortion and authentic distortion dataset in nature IQA also have different activation patterns.
+We observe that the activation patterns are different for different scenarios.
+Specifically, the natural IQA datasets, e.g., LIVE, CSIQ, KADID, show different activation patterns from the face IQA dataset GFIQA and the underwater IQA dataset UWIQA.
+The synthetic and authentic distortion datasets in natural IQA also have different activation patterns.
 These indicate that our MoAE module can assign experts with different activation levels to images of different scenarios, thereby capturing the discriminative features effectively.
@@
 They collaborate to achieve the powerful image assessment model Gamma.
-4.5 Generalization Capability Validation
+Figure 6. The change of merging factor $\sigma$ during training. The $\sigma$ is from the MoAE of the last visual encoder layer.
+4.5. Comparison with Mixed Training Methods
+In this subsection, we conduct a more detailed comparison with other mixed training methods.
+We first compare with LIQE and UNIQUE using the same training data and data splitting ratios.
+As shown in Table 7, our method achieves better performance on most datasets than LIQE and UNIQUE, especially on the KADID (+2.5% SRCC) and BID (+2.6% SRCC) datasets compared with LIQE.
+On other datasets, i.e., LIVE and LIVEC, our model also achieves competitive results.
+Overall, our model has superior performance on these five datasets.
+Compared with Q-Align, as shown in Table 8, our method achieves better results on KonIQ and KADID, and is also highly competitive on SPAQ.
+4.6. Generalization Capability Validation
 We further validate the generalization capability of our method on two datasets, exBeDDE and ECIQAD.
-The exBeDDE is a dehazed IQA dataset and the ECIQAD is an enhanced colonoscopy IQA dataset, which belong to completely different evaluation domains compared to the used datasets in mixed training.
+The exBeDDE is a dehazed IQA dataset and the ECIQAD is an enhanced colonoscopy IQA dataset, which belong to completely different domains compared to the used datasets in mixed training.
 We use naive prompt strategy for training and testing.
@@
 This indicates that our pretrained Gamma can be an effective foundation model to aid other assessment fields.
-5 Conclusion
+5. Conclusion
 This paper introduces Gamma, a generic image assessment model that can be applied to various image scenarios.
@@
 Extensive experiments demonstrate that our method can achieve SOTA performance on various datasets simultaneously, showing the strong generalization and general image assessment capabilities.
-Impact Statement
-This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
 Appendix A More Implementation Details
-A.1 Evaluation Criteria
-We use Spearman’s Rank-Order Correlation Coefficient (SRCC) and Pearson’s Linear Correlation Coefficient (PLCC) as criteria to measure the performance of IQA models.
-Both coefficients range from 0 to 1, with higher values indicating better performance.
-A.2 Training Details
-We follow the typical training strategy to fine-tune each dataset, including random cropping and random horizontal flipping.
-We conduct all experiments on 3090 GPU.
-Mixed training of the 12 datasets takes 10 hours on a 3090 GPU.
-For the task-specific training, Table 8 shows the detailed training setting for the different datasets.
-We use the learning rate of 2e-5 for all datasets.
-Table 8: Training settings for different datasets.
+Table 10. Training settings for different datasets.
 Dataset Task Epoch Batch size
 LIVE (Sheikh et al., 2006) SDN-IQA 50 8
-CSIQ (Larson & Chandler, 2010) SDN-IQA 50 8
+CSIQ (Larson and Chandler, 2010) SDN-IQA 50 8
 TID2013 (Ponomarenko et al., 2013) SDN-IQA 20 8
 KADID (Lin et al., 2019) SDN-IQA 20 8
-CLIVE (Ghadiyaram & Bovik, 2015) ADN-IQA 50 8
+CLIVE (Ghadiyaram and Bovik, 2015) ADN-IQA 50 8
 KonIQ (Hosu et al., 2020) ADN-IQA 20 8
 SPAQ (Fang et al., 2020) ADN-IQA 20 8
-GFIQA20k (Su et al., 2023b) F-IQA 10 8
-AGIQA3k (Li et al., 2023a) AG-IQA 20 8
+GFIQA20k (Su et al., 2023a) F-IQA 10 8
+AGIQA3k (Li et al., 2023b) AG-IQA 20 8
 UWIQA (Yang et al., 2021a) U-IQA 50 8
@@
 ECIQAD (Yue et al., 2023) EC-IQA 20 8
-A.3 Datasets
-In this paper, we use a total of 14 datasets, 12 of which are used for unified training and 2 are used to evaluate the generalization ability of our model.
-We present the details of the used datasets in Table 9.
-Table 9: Detail information about the 14 used datasets.
-Dataset Task Image Number Label Type Range
+Table 11. Detail information about the 14 used datasets.
+Dataset Task Image Number MOS Type Range
 LIVE (Sheikh et al., 2006) SDN-IQA 779 DMOS [1, 100]
-CSIQ (Larson & Chandler, 2010) 866 DMOS [0, 1]
+CSIQ (Larson and Chandler, 2010) 866 DMOS [0, 1]
 TID2013 (Ponomarenko et al., 2013) 3,000 MOS [0, 9]
@@
 SPAQ (Fang et al., 2020) ADN-IQA 11,125 MOS [0, 100]
-LIVEC (Ghadiyaram & Bovik, 2015) 1,162 MOS [1, 100]
+LIVEC (Ghadiyaram and Bovik, 2015) 1,162 MOS [1, 100]
 KonIQ-10K (Hosu et al., 2020) 10,073 MOS [0, 100]
-GFIQA20k (Su et al., 2023a) F-IQA 19,988 MOS [0, 1]
-AGIQA3k (Li et al., 2023a) AG-IQA 2,982 MOS [0, 1]
+GFIQA20k (Su et al., 2023b) F-IQA 19,988 MOS [0, 1]
+AGIQA3k (Li et al., 2023b) AG-IQA 2,982 MOS [0, 1]
 UWIQA (Yang et al., 2021a) U-IQA 890 MOS [0, 1]
@@
 ECIQAD (Yue et al., 2023) EC-IQA 2400 MOS [1, 9]
-Table 10: Detail information about the 14 used datasets.
-Method Trainable Parms FLOPs Inference time KonIQ SRCC KADID SRCC
-Q-Align (Wu et al., 2023) 8.2B (8200M) - 0.1s 0.938 0.934
-LIQE (Zhang et al., 2023) 151M 17.40G 0.02s 0.919 0.930
-Gamma 122.8M 28.45G 0.025s 0.939 0.962
-A.4 Details of the Scene-based Differential Prompt
-In the Scene-based Differential Prompt, we use different prompts for datasets from different scene.
-Specifically, we divide datasets into five categories, i.e., natural IQA, AI-generated IQA, underwater IQA, face IQA, natural IAA.
-We present the details in Table 11.
-Table 11: Text prompts used in Scene-based Differential Prompt.
+Table 12.
+Cross-dataset validation when using the same training data as LIQE (Zhang et al., 2023) and UNIQUE (Zhang et al., 2021).
+The subscripts “$s$” and “$r$” stand for models trained on KADID and KonIQ, respectively.
+Gamma uses the same 6 datasets as LIQE and UNIQUE for training and validation.
+Gamma†† uses 12 datasets (include TID2012 and SPAQ) for training and performs zero-shot cross dataset validation on other datasets.
+Gamma†† significantly enhances the zero-shot performance on other datasets, showing the effectiveness of mix-training via MoAE and SDP.
+Dataset TID2013 SPAQ AIGC2023 Average
+NIQE 0.314 0.578 - 0.446
+DBCNNs 0.686 0.412 0.730 0.609
+PaQ2PiQ 0.423 0.823 0.643 0.630
+MUSIQr 0.584 0.853 0.736 0.724
+UNIQUE 0.768 0.838 0.761 0.789
+LIQE 0.811 0.881 0.744 0.812
+Gamma 0.805 0.894 0.770 0.823
+Gamma†† - - 0.818 -
+Table 13. Generalization capability validation on the BAID (Yi et al., 2023) and ICAA17K (He et al., 2023) datasets. We use the pre-trained weight of Gamma for model initialization.
+(a) Results on the BAID dataset.
+Method SRCC PLCC
+NIMA (Talebi and Milanfar, 2018) 0.393 0.382
+MPada (Sheng et al., 2018) 0.437 0.425
+MLSP (Hosu et al., 2019) 0.441 0.430
+BIAA (Zhu et al., 2020) 0.389 0.376
+TANet (He et al., 2022) 0.453 0.437
+BAID (Yi et al., 2023) 0.473 0.467
+Ours 0.496 0.556
+(b) Results on the ICAA17K dataset.
+Method SRCC PLCC
+NIMA (Talebi and Milanfar, 2018) 0.809 0.815
+MPada (Sheng et al., 2018) 0.810 0.819
+MLSP (Hosu et al., 2019) 0.826 0.837
+BIAA (Zhu et al., 2020) 0.820 0.829
+MUSIQ (Ke et al., 2021) 0.835 0.841
+TANet (He et al., 2022) 0.829 0.836
+MaxViT (Tu et al., 2022) 0.855 0.874
+ICAA (He et al., 2023) 0.873 0.890
+Ours 0.877 0.893
+Table 14. Text prompts used in Scene-based Differential Prompt.
 Dataset Prompt
@@
 natural good-aesthetics image, natural perfect-aesthetics image}
-Appendix B Comparison with other mixed training methods
-Table 12: Comparison with LIQE (Zhang et al., 2023) and UNIQUE (Zhang et al., 2021) when using the same training data.
-Dataset LIVE CSIQ KADID BID LIVEC KonIQ Average
-Method SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
-UNIQUE 0.961 0.952 0.902 0.921 0.884 0.885 0.852 0.875 0.854 0.884 0.895 0.900 0.891 0.903
-LIQE 0.970 0.951 0.936 0.939 0.930 0.931 0.875 0.900 0.904 0.910 0.919 0.908 0.922 0.923
-Gamma 0.960 0.947 0.936 0.957 0.955 0.956 0.901 0.925 0.890 0.915 0.933 0.946 0.929 0.941
-Table 13: Cross-dataset validation when using the same training data as LIQE and UNIQUE.
-The subscripts “$s$” and “$r$” stand for models trained on KADID and KonIQ, respectively.
-Gamma uses the same 6 datasets as LIQE and UNIQUE for training and validation.
-Gamma†† uses 12 datasets (include TID2012 and SPAQ) for training and performs zero-shot cross dataset validation on other datasets.
-Gamma†† significantly enhances the zero-shot performance on other datasets, showing the effectiveness of mix-training via MoAE and SDP methods.
-Dataset TID2013 SPAQ AIGC2023 Average
-NIQE 0.314 0.578 - 0.446
-DBCNNs 0.686 0.412 0.730 0.609
-PaQ2PiQ 0.423 0.823 0.643 0.630
-MUSIQr 0.584 0.853 0.736 0.724
-UNIQUE 0.768 0.838 0.761 0.789
-LIQE 0.811 0.881 0.744 0.812
-Gamma 0.805 0.894 0.770 0.823
-Gamma†† - - 0.818 -
-In this subsection, we conduct a more detailed comparison with other mixed training methods.
-We first compare with LIQE and UNIQUE using the same training data and data splitting ratios.
-As shown in Table 12, our method achieves better performance on most datasets than LIQE and UNIQUE, especially on the KADID (+2.5% SRCC) and BID (+2.6% SRCC) datasets compared with LIQE.
-On other datasets, i.e., LIVE and LIVEC, our model also achieves competitive results.
-Overall, our model has superior performance on these five datasets.
-In addition, we conduct cross dataset validation under this setting.
-As shown in Table 13, our method achieves highly competitive results on TID2013 and SPAQ, demonstrating the strong generalization capability of our method.
-Note that using mix-training of more dataset can significantly enhance the zero-shot performance on other datasets, i.e., 0.770 SRCC to 0.818 SRCC on AIGC2023 dataset, demonstrating the effectiveness of mix-training via MoAE and SDP methods.
-Compared with Q-Align, as shown in Table 14, our method achieves better results on KonIQ and KADID, and is also highly competitive on SPAQ.
-Table 14: Comparison with Q-Align (Wu et al., 2023) when using the same training data.
-Dataset KonIQ SPAQ KADID Average
-Method SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
-Q-Align 0.938 0.945 0.931 0.933 0.934 0.935 0.934 0.938
-Gamma 0.940 0.950 0.928 0.932 0.962 0.964 0.943 0.949
-Appendix C Sensitivity analysis of prompt
-We analyze the sensitivity of prompts when the model is trained with scene-based differential prompts (SDP).
-Table 15 shows that using prompts different from SDP slightly reduces performance on most datasets, showing the robustness of our method.
-The quality prompt performs better than the general prompt on the IQA task, but performs worse on the IAA task, indicating the importance of appropriate prompts.
-In conclusion, our method is robust and insensitive to prompts, nevertheless we suggest using correct prompts to obtain better performance.
-Table 15: Sensitivity analysis of prompt.
+Table 15. Model complexity comparison with LIQE and Q-Align.
+Method Trainable Parms FLOPs Inference time KonIQ SRCC KADID SRCC
+Q-Align (Wu et al., 2023) 8.2B (8200M) - 0.1s 0.938 0.934
+LIQE (Zhang et al., 2023) 151M 17.40G 0.02s 0.919 0.930
+Gamma 122.8M 28.45G 0.025s 0.939 0.962
+Table 16.
+Sensitivity analysis of prompt.
 Quality prompt is {bad-quality, poor-quality, fair-quality, good-quality, perfect-quality}; General prompt replaces the scene prompt (see Table 14) to “general”, e.g., {underwear bad-quality image} to {general bad-quality image}.
@@
 SDP 0.891 0.914 0.939 0.950 0.953 0.953 0.960 0.968 0.887 0.923 0.873 0.884 0.750 0.749
+A.1. Evaluation Criteria
+We use Spearman’s Rank-Order Correlation Coefficient (SRCC) and Pearson’s Linear Correlation Coefficient (PLCC) as criteria to measure the performance of IQA models.
+Both coefficients range from 0 to 1, with higher values indicating better performance.
+We randomly crop a 224x224 image patch from the original image and perform score prediction, following (Qin et al., 2023; Zhang et al., 2023).
+This process is repeated 10 times, and the average of the 10 predicted scores is used as the final quality score of the image.
+A.2. Training Details
+We follow the typical training strategy to fine-tune each dataset, including random cropping and random horizontal flipping.
+We conduct all experiments on 3090 GPU.
+Mixed training of the 12 datasets takes 10 hours on a 3090 GPU.
+For the task-specific training, Table 10 shows the detailed training setting for the different datasets.
+We use the learning rate of 2e-5 for all datasets.
+A.3. Datasets
+In this paper, we use a total of 14 datasets, 12 of which are used for unified training and 2 are used to evaluate the generalization ability of our model.
+We present the details of the used datasets in Table 11.
+A.4. Details of the Scene-based Differential Prompt
+In the Scene-based Differential Prompt, we use different prompts for datasets from different scene.
+Specifically, we divide datasets into five categories, i.e., natural IQA, AI-generated IQA, underwater IQA, face IQA, natural IAA.
+We present the details in Table 14.
+Appendix B Cross Dataset Validation
+We conduct cross dataset validation to evaluate the generalization ability of our method.
+We compare with LIQE and UNIQUE using the same training data and data splitting ratios.
+As shown in Table 12, our method achieves highly competitive results on TID2013 and SPAQ, demonstrating the strong generalization capability of our method.
+Note that using mix-training of more dataset can significantly enhance the zero-shot performance on other datasets, i.e., 0.770 SRCC to 0.818 SRCC on AIGC2023 dataset, demonstrating the effectiveness of mix-training via MoAE and SDP methods.
+Appendix C More Generalization Capability Validation
+Gamma can be used as a pre-trained model and easily applied to new tasks after fine-tuning.
+We select the artistic IAA dataset BAID (Yi et al., 2023) and color IAA dataset ICAA17K (He et al., 2023) for further performance evaluation.
+BAID is a large-scale artistic image aesthetics assessment (AIAA) dataset, which consists of 60,337 artistic images covering various art forms.
+ICAA17K is a color-oriented IAA dataset, containing 17K images.
+We fine-tune Gamma for 10 epochs on them.
+As shown in the Table 13, Gamma achieves SOTA performance on two datasets, although it is not specifically designed for them.
+These results further verify the powerful generalization ability of Gamma.
 Appendix D Model efficiency analysis
@@
 Compared with Q-Align, we not only have better performance, but also have significantly lower model parameters and inference latency.
+Appendix E Sensitivity analysis of prompt
+We analyze the sensitivity of prompts when the model is trained with scene-based differential prompts (SDP).
+Table 16 shows that using prompts different from SDP slightly reduces performance on most datasets, showing the robustness of our method.
+The quality prompt performs better than the general prompt on the IQA task, but performs worse on the IAA task, indicating the importance of appropriate prompts.
+In conclusion, our method is robust and insensitive to prompts, nevertheless we suggest using correct prompts to obtain better performance.